accounting

1-Apple company purchases furniture’s at a cash price of $16,000. Related

expenditures are sales taxes $1,500, Insurance after the installation is $250,

Installation and testing $200. The furniture’s estimated useful life of is 5 years

and the salvage value is $1,500

 

a.       Determine the cost component of the furniture.

b.       Compute the total costs of the furniture.                 

c.       Using straight line method, how much will be the annual depreciation cost?

d.       What is the book value of the machine at the end of the 4th year? 

e.       What is the book value of the machine at the end of the 5th year?

 

 

2-XYZ  purchases factory machine at a cash price of $70,000 . Related

expenditures are sales taxes $2000, Insurance after the installation is $1,500,

Installation and testing $500, Salvage value is $500. Useful life of the machine is

5 years.

 

a.       Compute the component costs of the machine

b.       Compute the total costs of the machine

c.       Using straight line method, how much will be the annual depreciation cost?

d.   
Answer #1

1.

Given Data:

Totalcost of furniture = 16000 + 1500 + 250 +200 = $17950

Salvage value = $1500

Useful life = 5 years

Straight line Method of Depreciation

Depreciation cost = [Total cost - Salvage value] / Useful life

Depreciation cost = [17950 - 1500] / 5

Depreciation cost = $3290

Book value in 4th year = 17950 - 4 * 3290 = $4790

Book value in 5th year = $1500

2.

Given Data:

Totalcost of furniture = 70000 + 2000 + 1500 +500 = $74000

Salvage value = $500

Useful life = 5 years

Straight line Method of Depreciation

Depreciation cost = [Total cost - Salvage value] / Useful life

Depreciation cost = [74000- 500] / 5

Depreciation cost (Annual) = $14700